OpenStudy (anonymous):

This year, your brother Jack will be 2 years from being twice as old as your sister Jen. The sum of Jack's age and three times Jen's age is 66. How old is jen?

OpenStudy (anonymous):

(1) Jack-2=2*Jen or (2) Jack=(2*Jen)+2 (3) Jack+3*Jen=66 substituting (2) into (3) (4) 2*Jen+2+3*Jen=66 collecting terms (5) 5*Jen=66-2=64 (6) Jen=64/5=12.8

OpenStudy (anonymous):

it;s ok tat u type slow lyam. take ur time as long as u r trying hard! :)

OpenStudy (anonymous):

wait, so jen is 12.8 years old? does tat make sense?

OpenStudy (anonymous):

I think so, she is about 2 years and 9 months? Since Jack-2=2*Jen, and Jen=12.8 Jack=2*12.8+2=27.6

OpenStudy (anonymous):

lwt me double check my work

OpenStudy (anonymous):

ok cuz tat doesnt make much sense to me i think it should be whole numbers...just sayin

OpenStudy (anonymous):

Jack+3*Jen=66 27.6+3*12.8=66 so I think it is right

OpenStudy (anonymous):

o yeh, ur rite thx! again

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:) ur rly smart...a lifesaver

OpenStudy (anonymous):

fan me! I like ot help